in layman's term please answer these questions about Bismuth-213 (46 min) and cite at least 1 source in APA style

Assignment

Using your book and Internet resources, find five interesting pieces of information on the radioisotope you selected. Suggested pieces of information may include:

1. How is the radioisotope formed in nature (type of decay)?

2. Risks associated with the radioisotope

3. Number of subatomic nucleons present

4. Stability of the radioisotope

5. How this radioisotope is used.

6. Is the radioisotope used as a medical tracer? If yes, for what body part or process? Briefly explain.

this is what i have so far

-----------------------------------------------------------------------------------

213Bi is a mixed alpha and beta emitter having a half-life of 46 min. The primary mode of decay is by beta emission to the very short-lived, alpha emitter 213Po. Bismuth (213Bi) is formed from 225Ac decay. 225Ac is a pure alpha emitter with a half-life of 10 days.  

Both 212Bi and 213Bi used in medicine for radio-immunotherapy as a monoclonal antibodies to treat cancer cells from melanoma and ovarian cancer. It uses targeted alpha therapy (TAT), especially cancers, as it has a high energy (8.4 MeV).On the other hand, 212Bi is being studied for radio-immunotherapy of leukemia and for targeting the vascular endothelial cells.
Answer #1

1. The radioisotope is Bismuth- 213 (213 Bi)

(a) The isotope is formed by beta emission of 213Po:

213Po84 - 0 e -1 = 213 Bi83

(b) It is also produced by three consecutive alpha emissions from 225Ac:

225Ac89 - 4He2 = 221Fr87

221Fr87 - 4He2 = 217At85

217At85 - 4He2 = 213 Bi83

2. This radioisotpe is alpha emitter and alpha rays are very harmful for living cells and organs.

3. Nucleons are neutrons and protons in the nucleus.

Number of protons in the radioisotope is 83

Number of protons is (213-83) = 130

4. Being a radioisotope it is unstable with halflife of 46 minutes. That means half amount of the isotope decays in 46 minutes. It emits alpha ray to produce 209Tl

5. 213Bi used in medicine for radio-immunotherapy as a monoclonal antibodies to treat cancer cells from melanoma and ovarian cancer. It uses targeted alpha therapy (TAT), especially cancers, as it has a high energy (8.4 MeV).